OpenStudy (anonymous):

What is the equation of the line that passes through (12, 4) and is perpendicular to the graph of y = –3/4x – 2?

OpenStudy (anonymous):

y = –3/4x – 2? a) y = 4/3 x – 12 ....................Ans b) y = 3/4x – 5 c) y = –4/3x + 20 d) y = –3/4x + 13

OpenStudy (anonymous):

Just by looking at the answers' slope values you can get this one. Perp lines have slopes that are the negative reciprocal of each other. so 4/3 should be the slope....